OpenStudy (anonymous):

solve 3x - 9 + x > 27

OpenStudy (astrophysics):

Start of by collecting like terms on the left side, \[3x+x = 4x \implies 4x-9 > 2\] now add 9 on both sides to get rid of it on the left, \[4x>2+9 \implies 4x> 11\] divide by 4, \[x > \frac{ 11 }{ 4 }\]

OpenStudy (anonymous):

Treat the > like an =. The goal is to isolate x, so I would add 9 to both sides of the equation to start out.

OpenStudy (anonymous):

that answer isnt a choice

OpenStudy (anonymous):

This will give you 3x+x=36 Combine like terms to get 4x=36. Divide 36 by 4 to isolate x, so x=39.

OpenStudy (astrophysics):

Oh I see the problem I put >2 not 27

OpenStudy (astrophysics):

XD

OpenStudy (anonymous):

oops. x=9

OpenStudy (astrophysics):

x >9

OpenStudy (anonymous):

thank you guys so much

OpenStudy (anonymous):

Your welcome.